OpenStudy (anonymous):

Midtown Delivery Service delivers packages which cost $2.40 per package to deliver. The fixed cost to run the delivery truck is $376 per day. If the company charges $6.40 per package, how many packages must be delivered daily to make a profit of

OpenStudy (anonymous):

Let p be the number of packages delivered.\[\text{OneDayProfit}=\frac{64}{10}p-\left(\frac{24}{10}p+376\right)=4 p-376 \]Set the profit expression to zero and solve for p, the number or packages required for zero profit.\[4 p-376=0 \]\[p=94 \]At least 95 packages must be delivered to make a profit.

OpenStudy (anonymous):

Thank you for the medal. There is a spelling error in the solution text. The word "or" should have been typed "of".